40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 19 18:41:42 2022, yuta, Update, BHD, Contrast measurements for Michelson and ITM-LO ContrastMeasurements.pdf
    Reply  Wed Jul 20 18:07:52 2022, Paco, Update, BHD, BHD MICH test BHD_MICH_OSC.pdf
       Reply  Tue Jul 26 20:54:08 2022, Paco, Update, BHD, BHD MICH test - LO phase control 20220726_BICHD.pdf
          Reply  Wed Jul 27 18:30:50 2022, yuta, Update, BHD, LO beam power at BHD DCPDs is significantly lower than expected LOBeamAtBHD.JPGITMXSingleBounceAtBHD.jpg
             Reply  Thu Jul 28 14:34:40 2022, Yehonathan, Update, BHD, LO beam power at BHD DCPDs is significantly lower than expected 
                Reply  Thu Jul 28 16:51:55 2022, Tega, Update, BHD, Shaking test for LO beam AS beam to BHD DCPDs Screenshot_2022-07-28_16-58-34_LOandASShaking.pngScreenshot_2022-07-28_18-08-55_DCPDPOPSuspensionCoherence.png
                   Reply  Thu Jul 28 20:16:26 2022, Anchal, Update, BHD, Shaking test for LO beam AS beam to BHD DCPDs 
                   Reply  Fri Jul 29 18:24:53 2022, Tega, Update, BHD, LO beam power improved by factor of 6 after LO and AS beam alignment LO_and_oldAS_settings.pngBHD_fringe_settings.pngScreenshot_2022-07-29_19-37-39.pngFromTheLeft-AS-POP-LO.JPG
                      Reply  Wed Aug 3 16:00:51 2022, yuta, Update, BHD, BHD fringe aligned with reduced LO and AS beam clipping Screenshot_2022-08-03_15-46-36_BHDfringeAlmostUnclipped.png
                         Reply  Tue Aug 9 15:33:12 2022, yuta, Update, BHD, BHD fringe contrast improved from 43% to 74% HPC-DCPD_B_OUT_1344118517_ITMY-LO.pngHPC-DCPD_A_OUT_1344118517_ITMY-LO.pngHPC-DCPD_B_OUT_1344118318_ITMX-LO.pngHPC-DCPD_A_OUT_1344118318_ITMX-LO.png
                            Reply  Tue Aug 9 15:50:22 2022, Koji, Update, BHD, BHD fringe contrast improved from 43% to 74% 
                               Reply  Wed Nov 16 15:09:08 2022, yuta, Update, BHD, BHD fringe contrast measured with unwhitening filters ContrastMeasurements20221116_edited.pdf
                         Reply  Tue Oct 25 08:25:00 2022, JC, Update, BHD, BHD fringe aligned with reduced LO and AS beam clipping 
                         Reply  Thu Nov 17 11:24:39 2022, JC, HowTo, LSC, Locking MICH 
                            Reply  Tue Nov 29 11:38:37 2022, JC, HowTo, LSC, Lock Single Arm After MICH lock Screen_Shot_2022-11-29_at_11.01.56_AM.png
Message ID: 17037     Entry time: Tue Jul 26 20:54:08 2022     In reply to: 17024     Reply to this: 17040
Author: Paco 
Type: Update 
Category: BHD 
Subject: BHD MICH test - LO phase control 

[Yuta, Paco]


TL;DR Successfully controlled LO phase, and did BHD-MICH readout with various MICH offsets and LO phases.


Today we implemented a DCPD based LO phase control. First, we remeasured the balancing gain at 311.1 Hz (the MICH oscillator freq) and combined C1:HPC-DCPD_A_OUT with C1:HPC-DCPD_B_OUT to produce the balanced homodyne error signal (A-B). We feed this error signal to C1:HPC-LO_PHASE_IN1 and for the main loop filters we simply recycled the LSC-MICH loop filters FM2 through FM5 (we also copied FM8, but didn't end up using it much). Then, we verified the LO phase can be controlled by actuating either on LO1 or LO2. For LO2, we added an oscillator in the HPC LOCKINS at 318.75 Hz (we kept this on at 1000 counts for the measurements below).

The LO phase control was achieved with a loop gain in the range of 10-30 (we used 20), no offset, and FM4, and FM5 engaged. FM2 can be added to boost, but we usually skipped FM3. Then, we went through a set of measurements similar to the ones described in a previous elog. A key difference with respect to the measurements from before is that we locked MICH using AS55Q (as opposed to REFL55Q). This allowed us to reach higher MICH offsets without losing lock. After turning on the MICH oscillator at 3000 counts, we looked at:

  1. LO misaligned + MICH at dark fringe (offset = -21).
    • Here, we don't expect to see any MICH signal and indeed we don't, except for a small residual peak from perhaps a MICH offset or slightly imbalanced PDs.
  2. LO aligned, but uncontrolled + MICH at dark fringe (offset = -21).
    • Here we would naively expect MICH to show up in A-B, but because of the uncontrolled LO phase, we mostly see the noise baseline (mostly from LO RIN? ...see measurement 3) under which this signal is probably buried. Indeed, the LO fringe increased noise in A, B, and A-B but not in A+B. This is nice. yes
  3. LO aligned, but uncontrolled + MICH with dc readout (offset = +50).
    • Here we expected the MICH signal to show up due to the large offset, and we can indeed see it in A, B, and A+B, but not in A-B. Nevertheless we see almost exactly the same noise level even though we allow some AS light into the BHD readout, so maybe the noise observed in the A-B channel from measurements 2 and 3 is mostly from LO RIN. This needs further investigation...
  4. LO aligned, controlled at no offset + MICH with dc readout (offset = +50).
    • In general here we expected to see a noise reduction in the A-B channel since the LO fringe is stable, and a MICH signal should appear. Furthermore, since LO phase is under control, we expect the LO2 Oscillator to appear which it does for this and the following measurements. Because of the relative freedom, we tried this measurement in two cases:
      1. When feeding back to LO1
        • We actually see MICH in the A-B channel, as expected, after the noise level dropped by ~ 5. We also observed small sidebands +- 1 Hz away from the MICH peak, probably due to local damping in either LO or AS paths.
      2. When feeding back to LO2
        • We also see MICH here, with a slightly better drop in noise (relative to feeding back to LO1). Sidebands persisted here, but around at +- 2 Hz.
  5. LO aligned, controlled (offset = 10) + MICH with dc readout (offset = +50). *
    • Here, we expected the A-B MICH content to increase dramatically, and indeed it does after a little tuning of the LO phase heart. The noise level decreased slightly because LO phase noise is decreased around the optimal point.
  6. LO aligned, controlled (offset = 20) + MICH with dc readout (offset = +30). *
    • Here, we naively expected A+B MICH content to decrease, but A-B remain constant. In order to see this we tried to keep the balance between the offsets, but this was hard. We don't really see much of this effect, so this also needs further investigation. As long as we keep controlling the LO phase using the DCPDs because the offsets tend to reduce the error signal we will have a harder time.

* For these measurements we actuated on LO2 to keep the LO phase under control.

Note that the color code above corresponds to the traces shown in Attachment #1.


What's next?

  • Alignment of LO and AS might be far from optimized, so it should be tried more seriously.
  • What's the actual LO power? How does it compare with AS power at whatever MICH offsets?
  • Try audio dither LO phase control.
    • With MICH offset.
    • Without MICH offset, double demod (after dolphin fix crying)
Attachment 1: 20220726_BICHD.pdf  28 kB  Uploaded Tue Jul 26 22:20:51 2022  | Hide | Hide all
20220726_BICHD.pdf
ELOG V3.1.3-